# New ultralight scalar particles and the mass-radius relation of white dwarfs - the important role of Sirius B

### v1

@@ 4 Mass-radius relation of white dwarfs
In the next step, we have calculated white dwarf models and mass-radius relations with different versions of our non-classical EOS leading to first-order phase transitions (FOPT). Fig. 5 shows the result for the Scalar EOS with coupling constant $c_{m}=0.03$. Obviously, the models fail badly to reproduce the observed MRR. The metastable models lead to better results but their branches stop at $\sim$0.55 M.
Fig. 6 investigates the dependence of the non-classical FOPT EOS on the coupling parameter $c_{m}$. It shows the MRR at Teff = 25000 K with the parameter $c_{m}$ in the range 0.01 to 0.14. All models with a phase transition fail badly to reproduce the observed radii. For $c_{m}\gtrsim 0.03$, the sourced branch of the EOS is in tension with all considered white dwarfs. However, the unsourced and metastable branches of the EOS come closer to the observations. Increasing the value of $c_{m}$ increases the critical density (see Eq. (6) and Fig. 2) and thus the fraction of white dwarfs that can be explained even in the presence of a scalar field. But, very obviously, even the model with the largest possible $c_{m}$, $c_{m}=0.14$, cannot reproduce the radius of Sirius B. On the other hand, models with smaller $c_{m}$ around $c_{m}\approx 0.01$ fit Sirius B nicely, but do not reproduce the radii of the less massive objects. This strong deviation with respect to the other models can be understood from the EOS (compare Fig. 2, bottom), which shows a corresponding strong deviation through a large part of the white dwarf stellar interior. At this small value of $c_{m}$, the phase transition is very strongly first-order since it is very close to the case of an NGS. To sum up, the MRR provides a crucial constraint on the nature of these particles and rules out the complete range of coupling parameters leading to an FOPT.
@@ 5 Results and Conclusions
For the scalar field coupling to electron the comparison with the observed white dwarf mass radius data has provided a very strong constraint. All models that lead to a first-order phase transition can be excluded. The part of the theoretical MR-relations corresponding to the metastable branch of the EOS can reproduce some of the observed data at lower masses, in particular, when $c_{m}$ (or the critical density $\rho_{c}$) is increased. However, for the highest possible value leading to an FOPT $c_{m}$ = 0.14, the resulting mass and radius is in clear disagreement with the observed values of Sirius B. Given the precision of the determination of the mass and radius of Sirius B, this is a very clear exclusion.
@@
Above and to the left of the red area, $c_{m}$ becomes lower. At $c_{m}\lesssim 0.0093$, the scalar field no longer leads to a first-order phase transition but a new ground state in the white dwarf EOS. The corresponding blue-grey area in Fig. 8 has been investigated by (Balkin et al., 2024, 2025) and (Bartnick et al., 2025): The new ground state EOS leads to a complete change of the structure of white dwarfs and to forbidden gaps in their mass radius relation in clear disagreement with the observations.
For lower critical densities, the new ground state would also occur in the Sun. Through a considerably lower electron mass in the solar atmosphere, this would lead to solar spectra incompatible with observations (Bartnick et al., 2025). This area is colored yellow. The grey exclusion area results from the observation of rapidly rotating solar mass black holes, which would not exist with light scalars in this parameter domain as they extract energy from the black hole through "superradiance" leading to a spin-down (Brito et al., 2015; Witte & Mummery, 2025). The blue exclusion area results from precision experiments on Earth using quantum clocks and interferometry (Kennedy et al., 2020; Oswald et al., 2022; Vermeulen et al., 2021; Savalle et al., 2021; Aharony et al., 2021; Antypas et al., 2019; Aiello et al., 2022; Branca et al., 2017; Sherrill et al., 2023; Kobayashi et al., 2022; Campbell et al., 2021; Zhang et al., 2023; Oswald et al., 2025). Note that these experimental bounds only apply if the scalar field accounts for a large fraction of the dark matter abundance. Here, we translate the bounds directly from the linear case, assuming the background dark matter density. There might be changes due to finite density effects leading to non-trivial profiles around the Earth or the Sun (similar to sourcing, but without the scalar mass becoming tachyonic), see (Hees et al., 2018; Banerjee et al., 2023; Budker et al., 2023; Banerjee et al., 2025b; del Castillo et al., 2025; Grossman et al., 2025; Gan et al., 2025).
@@
Figure 9: Constraints on $\mathbb{Z}_{\mathcal{N}}$ axion models as a function of axion mass $m_{a}$ and decay constant $f_{a}$. Constraints due to a (modified) crossover phase transition in white dwarfs derived in this work are shown in red. The $m_{a}-f_{a}$-relation in the case $\mathcal{N}=1$, the standard QCD axion, is given by the yellow line; other constraints are discussed in the main text. Plot modified from (O’Hare, 2020).
@@
The white area below the new red exclusion area corresponds to smaller $\mathcal{N}$. Here, phase transitions occur in a similar way as for $\mathcal{N}=31$ but the corresponding metastable branch is able to reproduce the radius of Sirius and no exclusion is obtained. As in the scalar case, if the decay constant is too large (i.e. weak coupling limit), gradient effects prevent axion sourcing and thus the crossover phase transition in white dwarfs, which ends the exclusion for large $f_{a}$. In principle, the precise shape of the exclusion at large $f_{a}$ requires solving the full coupled equations describing the scalar field and the white dwarf, which has been done in (Balkin et al., 2024, 2025) but goes beyond the scope of this work. Consequently, as in the scalar case discussed above, we set the limit by estimating the typical scale of the axion gradient and comparing it to the white dwarf size. To be conservative, we end the exclusion slightly earlier compared to (Balkin et al., 2024), where full gradient effects are computed.
@@
The exclusion areas at higher coupling constants (large $\mathcal{N}$) results from a new ground state (NGS) in white dwarfs (Balkin et al., 2024) and sourcing in the solar core (Hook & Huang, 2018). The bottom green area is also related to axion sourcing, in particular the influence it would have on the neutron star merger GW170817 (Zhang et al., 2021).
@@
Again, black hole superradiance (Witte & Mummery, 2025) leads to the exclusion shown in grey. Note that, for the $\mathbb{Z}_{\mathcal{N}}$-axion, the quartic-self interactions $\lambda a^{4}$ are enhanced by a factor $\mathcal{N}^{2}$ at fixed mass and decay constant compared to the normal QCD-axion case; consequently, the superradiance bounds are correspondingly weaker. Axion cooling leads to the constraints obtained from supernova 1987A (Springmann et al., 2024). Finally, experiments such as Beam EDM (Schulthess et al., 2022) lead to constraints at very low axion mass, if an axion dark matter background is assumed (Banerjee et al., 2025b; del Castillo et al., 2025). We finally note that sourcing in neutron stars can also constrain light axions due to anomalously fast cooling (Gómez-Bañón et al., 2024; Kumamoto et al., 2025). Since these limits rely on the presence of a new ground state for which heat-blanketing envelopes can be absent, they are fundamentally different from traditional cooling bounds (such as Raffelt (1996); Buschmann et al. (2022); Springmann et al. (2025)). Importantly, the $\mathbb{Z}_{\mathcal{N}}$-potential has no new ground state at neutron star densities (Balkin et al., 2024) and hence avoids these limits, which is why they are omitted in Fig. 9.
@@ Acknowledgements
The authors (KB, RK, AW) acknowledge support by the Munich Excellence Cluster Origins and the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2094 390783311. KB and AW are also supported by the Collaborative Research Center SFB1258. SS is supported by the Swiss National Science Foundation under contract 200020-213104 and acknowledges the hospitality of the CERN theory group. KS is supported by a research grant from Mr. and Mrs. George Zbeda and by the Minerva Foundation.
@@ A.1 Scalar field coupling to electrons
Here, we only include quadratic couplings between the scalar field and the fermions. Since this Appendix focuses on particle physics details, we will use natural units where $\hbar=c=1$ from now on. In general, linear couplings would also be present, but they are tightly constrained experimentally, e.g., from fifth force searches (see e.g. (Schlamminger et al., 2008; Lee et al., 2020; Tan et al., 2020)). By imposing a discrete symmetry $\mathbb{Z}_{2}$ symmetry $\phi\to-\phi$, linear couplings are forbidden, making the leading interactions quadratic in $\phi$. In general, $d_{m_{e}}^{(2)}$ could have any sign; here, we focus on the negative case $d_{m_{e}}^{(2)}<0$, which has profound implications on stellar structure.
@@
Deriving the full in density scalar equations of motion, one finds a coupled system together with the equations of stellar structure, describing stars with sourced scalar fields, see (Balkin et al., 2024, 2025) for the limit of negligible temperature. Solving this system in full generality goes beyond the scope of this work. In most of the parameter space a simplifying limit exists. Gradients of the scalar field $\phi$ have an associated energy cost, and the field will change on scales of the order of
@@
a potential barrier prevents the scalar field from moving away from 0. Here $V_{\mathrm{eff}}(\phi,n_{e})=1/2\,m_{\phi}^{2}\phi^{2}+\varepsilon_{e}(n_{e},m_{e}(\phi))$ is the effective potential at constant number density. Its minimum at a given number density (so again the solution of the scalar equation of motion) also defines $\phi(n)$ as discussed in the main text. In general, the critical density $n_{c}$ is larger than the density where $\phi_{\mathrm{eq}}=\phi(\mu)$ starts to deviate from 0. Tunneling into the equilibrium state is extremely suppressed due to the large field excursions, so if central density of a star is below the critical one, the field will remain at 0. This is analogous to supercooling in an ordinary phase transition. On the other hand, once the critical density is surpassed somewhere within the star, $\phi$ will indeed globally track its equilibrium solution.
@@ A.2 Lighter version of the QCD axion
Another well-motivated example for a light scalar field quadratically coupled to the Standard Model is the QCD axion $a$. The defining $aG\tilde{G}$ coupling, where $G$ and $\tilde{G}$ are the gluon field strength and its dual, necessary to solve the strong CP-problem, at low energies, leads to a quadratic axion-nucleon coupling of the form (Hook & Huang, 2018; Springmann et al., 2025)
@@
Here $N=(n,p)$ is the nucleon-doublet, $f_{a}$ the axion-decay constant, $z=m_{u}/m_{d}$ is the up to down quark mass ratio, and $\sigma_{N}\simeq 50~\mathrm{MeV}$ the nucleon sigma term. In the same way as described in the case of a scalar field coupling to electrons, this coupling can source the axion in dense matter as shown in (Hook & Huang, 2018; Balkin et al., 2024, 2025). For a normal QCD-axion, where the height of the potential, or equivalently, the mass at given $f_{a}$, is fully fixed, this would only occur above nuclear saturation density, where one lacks perturbative control. If the height of the potential was lower or the mass of the axion lighter at a given $f_{a}$, sourcing occurs at lower densities and can be probed and constrained in white dwarfs (Balkin et al., 2024). In particular, this can be achieved in a technically natural way with $\mathcal{N}$ copies of the Standard Model and enforcing a $\mathbb{Z}_{\mathcal{N}}$- symmetry as first described by (Hook, 2018). Requiring the strong CP-problem to still be solved fixes $\mathcal{N}$ to be odd. For large $\mathcal{N}$, the axion potential than can be approximated as (Di Luzio et al., 2021)
@@
where $m_{\pi}$ and $f_{\pi}$ are the pion mass and decay constant, respectively. In particular, the axion mass (at a given $f_{a}$) and thus the critical density are suppressed by a factor of roughly $z^{\mathcal{N}}\simeq(1/2)^{\mathcal{N}}$ with respect to the standard QCD-axion. As shown in (Balkin et al., 2024), values of $\mathcal{N}\geq 33$ lead to a new ground state of matter in white dwarfs and come with a large gap in the predicted radii of white dwarfs, which is incompatible with observations. For $\mathcal{N}=29$ and larger, the critical density is high enough for the metastable branch to explain all observed white dwarfs, and no exclusion can be placed. This leaves $\mathcal{N}=31$, which we focus on in this work.
@@
From the behavior of the potential, we can now understand the behavior of the field values and consequently the EOS. Starting from the unsourced case, the axion field sits at 0. Due to the potential barrier, the axion field remains in the metastable minimum at 0 until it fully vanishes at the critical density. This leads to a metastable branch of the EOS and is shown in red in Fig. 10. At densities above the critical density, the axion field can simply track the global minimum of the finite density effective potential at $a=\pi f_{a}$. If the axion has been sourced once within a star, it will continue to track the global minimum of the potential even at densities below $n_{c}$. For densities below the critical density, this global minimum at first stays at $\pi f_{a}$ and then continuously shifts toward the zero-density minimum closest to $\pi f_{a}$, at $\pi\left(1-\frac{1}{\mathcal{N}}\right)f_{a}$ as marked by the arrows in Fig. 10. Only at (near) zero-density, when the energy gain from the potential becomes lower than the gradient energy required for the transition to the true minimum, does the field return to the outside value of $a=0$.

### v2

@@ 4 Mass-radius relation of white dwarfs
In the next step, we have calculated white dwarf models and mass-radius relations with different versions of our non-classical EOS leading to first-order phase transitions. Fig. 5 shows the result for the scalar EOS with coupling constant $c_{m}=0.03$. Obviously, the models fail badly to reproduce the observed MRR. The metastable models lead to better results but their branches stop at $\sim$0.55 M.
Fig. 6 investigates the dependence of the non-classical FOPT EOS on the coupling parameter $c_{m}$. It shows the MRR at $T_{\mathrm{eff}}$ = 25 000 K with the parameter $c_{m}$ in the range 0.01 to 0.14. All models with a phase transition fail badly to reproduce the observed radii. For $c_{m}\gtrsim 0.03$, the sourced branch of the EOS is in tension with all considered white dwarfs. However, the unsourced and metastable branches of the EOS come closer to the observations. Increasing the value of $c_{m}$ increases the critical density (see Eq. (6) and Fig. 2) and thus the fraction of white dwarfs that can be explained even in the presence of a scalar field. But, very obviously, even the model with the largest possible $c_{m}$, $c_{m}=0.14$, cannot reproduce the radius of Sirius B. On the other hand, models with smaller $c_{m}$ around $c_{m}\approx 0.01$ fit Sirius B nicely, but do not reproduce the radii of the less massive objects. This strong deviation with respect to the other models can be understood from the EOS (compare Fig. 2, bottom), which shows a corresponding strong deviation through a large part of the white dwarf stellar interior. At this small value of $c_{m}$, the phase transition is very strongly first-order since it is very close to the case of an NGS. To sum up, the MRR provides a crucial constraint on the nature of these particles and rules out the complete range of coupling parameters leading to an FOPT.
@@ 5 Results and Conclusions
For the scalar field coupling to electron the comparison with the observed white dwarf mass radius data has provided a very strong constraint. All models that lead to a first-order phase transition can be excluded. The part of the theoretical MRRs corresponding to the metastable branch of the EOS can reproduce some of the observed data at lower masses, in particular, when $c_{m}$ (or the critical density $\rho_{c}$) is increased. However, for the highest possible value leading to an FOPT $c_{m}$ = 0.14, the resulting mass and radius is in clear disagreement with the observed values of Sirius B. Given the precision of the determination of the mass and radius of Sirius B, this is a very clear exclusion.
@@
Above and to the left of the red area, $c_{m}$ becomes lower. At $c_{m}\lesssim 0.0093$, the scalar field no longer leads to a first-order phase transition but a new ground state in the white dwarf EOS. The corresponding blue-grey area in Fig. 8 has been investigated by Balkin et al. (2024, 2025) and Bartnick et al. (2025): the new ground state EOS leads to a complete change of the structure of white dwarfs and to forbidden gaps in their mass radius relation in clear disagreement with the observations.
For lower critical densities, the new ground state would also occur in the Sun. Through a considerably lower electron mass in the solar atmosphere, this would lead to solar spectra incompatible with observations (Bartnick et al., 2025). This area is coloured yellow. The grey exclusion area results from the observation of rapidly rotating solar mass black holes, which would not exist with light scalars in this parameter domain as they extract energy from the black hole through "superradiance" leading to a spin-down (Brito et al., 2015; Witte & Mummery, 2025). The blue exclusion area results from precision experiments on Earth using quantum clocks and interferometry (Kennedy et al., 2020; Oswald et al., 2022; Vermeulen et al., 2021; Savalle et al., 2021; Aharony et al., 2021; Antypas et al., 2019; Aiello et al., 2022; Branca et al., 2017; Sherrill et al., 2023; Kobayashi et al., 2022; Campbell et al., 2021; Zhang et al., 2023; Oswald et al., 2025). Note that these experimental bounds only apply if the scalar field accounts for a large fraction of the dark matter abundance. Here, we translate the bounds directly from the linear case, assuming the background dark matter density. There might be changes due to finite density effects leading to non-trivial profiles around the Earth or the Sun (similar to sourcing, but without the scalar mass becoming tachyonic), see (Hees et al., 2018; Banerjee et al., 2023; Budker et al., 2023; Banerjee et al., 2025a; del Castillo et al., 2025; Grossman et al., 2025; Gan et al., 2025).
@@
Figure 9: Constraints on $\mathbb{Z}_{\mathcal{N}}$ axion models as a function of axion mass $m_{a}$ and decay constant $f_{a}$. Constraints due to a (modified) crossover phase transition in white dwarfs derived in this work are shown in red. The $m_{a}-f_{a}$ relation in the case $\mathcal{N}=1$, the standard QCD axion, is given by the yellow line; other constraints are discussed in the main text. Plot modified from O’Hare (2020).
@@
The white area below the new red exclusion area corresponds to smaller $\mathcal{N}$. Here, phase transitions occur in a similar way as for $\mathcal{N}=31$ but the corresponding metastable branch is able to reproduce the radius of Sirius and no exclusion is obtained. As in the scalar case, if the decay constant is too large (i.e. weak coupling limit), gradient effects prevent axion sourcing and thus the crossover phase transition in white dwarfs, which ends the exclusion for large $f_{a}$. In principle, the precise shape of the exclusion at large $f_{a}$ requires solving the full coupled equations describing the scalar field and the white dwarf, which has been done in (Balkin et al., 2024, 2025) but goes beyond the scope of this work. Consequently, as in the scalar case discussed above, we set the limit by estimating the typical scale of the axion gradient and comparing it to the white dwarf size. To be conservative, we end the exclusion slightly earlier compared to Balkin et al. (2024), where full gradient effects are computed.
@@
The exclusion areas at higher coupling constants (large $\mathcal{N}$) results from a NGS in white dwarfs (Balkin et al., 2024) and sourcing in the solar core (Hook & Huang, 2018). The bottom green area is also related to axion sourcing, in particular the influence it would have on the neutron star merger GW170817 (Zhang et al., 2021).
@@
Again, black hole superradiance (Witte & Mummery, 2025) leads to the exclusion shown in grey. Note that, for the $\mathbb{Z}_{\mathcal{N}}$-axion, the quartic-self interactions $\lambda a^{4}$ are enhanced by a factor $\mathcal{N}^{2}$ at fixed mass and decay constant compared to the normal QCD-axion case; consequently, the superradiance bounds are correspondingly weaker. Axion cooling leads to the constraints obtained from supernova 1987A (Springmann et al., 2025b). Finally, experiments such as Beam EDM (Schulthess et al., 2022) lead to constraints at very low axion mass, if an axion dark matter background is assumed (Banerjee et al., 2025a; del Castillo et al., 2025). We finally note that sourcing in neutron stars can also constrain light axions due to anomalously fast cooling (Gómez-Bañón et al., 2024; Kumamoto et al., 2025). Since these limits rely on the presence of a new ground state for which heat-blanketing envelopes can be absent, they are fundamentally different from traditional cooling bounds (such as Raffelt (1996); Buschmann et al. (2022); Springmann et al. (2025a)). Importantly, the $\mathbb{Z}_{\mathcal{N}}$-potential has no new ground state at neutron star densities (Balkin et al., 2024) and hence avoids these limits, which is why they are omitted in Fig. 9.
@@ Acknowledgements
The authors (KB, RK, AW) acknowledge support by the Munich Excellence Cluster Origins and the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311. KB and AW are also supported by the Collaborative Research Center SFB1258. SS is supported by the Swiss National Science Foundation under contract 200020-213104 and acknowledges the hospitality of the CERN theory group. KS is supported by a research grant from Mr. and Mrs. George Zbeda and by the Minerva Foundation.
@@ A.1 Scalar field coupling to electrons
Here, we only include quadratic couplings between the scalar field and the fermions. Since this Appendix focuses on particle physics details, we will use natural units where $\hbar=c=1$ from now on. In general, linear couplings would also be present, but they are tightly constrained experimentally, e.g., from fifth force searches (see e.g. Schlamminger et al., 2008; Lee et al., 2020; Tan et al., 2020). By imposing a discrete symmetry $\mathbb{Z}_{2}$ symmetry $\phi\to-\phi$, linear couplings are forbidden, making the leading interactions quadratic in $\phi$. In general, $d_{m_{e}}^{(2)}$ could have any sign; here, we focus on the negative case $d_{m_{e}}^{(2)}<0$, which has profound implications on stellar structure.
@@
Deriving the full in-density scalar equations of motion, one finds a coupled system together with the equations of stellar structure, describing stars with sourced scalar fields, see Balkin et al. (2024, 2025) for the limit of negligible temperature. Solving this system in full generality goes beyond the scope of this work. In most of the parameter space a simplifying limit exists. Gradients of the scalar field $\phi$ have an associated energy cost, and the field will change on scales of the order of
@@
a potential barrier prevents the scalar field from moving away from 0. Here $V_{\mathrm{eff}}(\phi,n_{e})=1/2\,m_{\phi}^{2}\phi^{2}+\varepsilon_{e}(n_{e},m_{e}(\phi))$ is the effective potential at constant number density. Its minimum at a given number density (so again the solution of the scalar equation of motion) also defines $\phi(n)$ as discussed in the main text. In general, the critical density $n_{c}$ is larger than the density where $\phi_{\mathrm{eq}}=\phi(\mu)$ starts to deviate from 0. Tunnelling into the equilibrium state is extremely suppressed due to the large field excursions, so if central density of a star is below the critical one, the field will remain at 0. This is analogous to supercooling in an ordinary phase transition. On the other hand, once the critical density is surpassed somewhere within the star, $\phi$ will indeed globally track its equilibrium solution.
@@ A.2 Lighter version of the QCD axion
Another well-motivated example for a light scalar field quadratically coupled to the Standard Model is the QCD axion $a$. The defining $aG\tilde{G}$ coupling, where $G$ and $\tilde{G}$ are the gluon field strength and its dual, necessary to solve the strong CP-problem, at low energies, leads to a quadratic axion-nucleon coupling of the form (Hook & Huang, 2018; Springmann et al., 2025a)
@@
Here $N=(n,p)$ is the nucleon-doublet, $f_{a}$ the axion-decay constant, $z=m_{u}/m_{d}$ is the up to down quark mass ratio, and $\sigma_{N}\simeq 50~\mathrm{MeV}$ the nucleon sigma term. In the same way as described in the case of a scalar field coupling to electrons, this coupling can source the axion in dense matter as shown in Hook & Huang (2018); Balkin et al. (2024, 2025). For a normal QCD-axion, where the height of the potential, or equivalently, the mass at given $f_{a}$, is fully fixed, this would only occur above nuclear saturation density, where one lacks perturbative control. If the height of the potential was lower or the mass of the axion lighter at a given $f_{a}$, sourcing occurs at lower densities and can be probed and constrained in white dwarfs (Balkin et al., 2024). In particular, this can be achieved in a technically natural way with $\mathcal{N}$ copies of the Standard Model and enforcing a $\mathbb{Z}_{\mathcal{N}}$- symmetry as first described by Hook (2018). Requiring the strong CP-problem to still be solved fixes $\mathcal{N}$ to be odd. For large $\mathcal{N}$, the axion potential than can be approximated as (Di Luzio et al., 2021)
@@
where $m_{\pi}$ and $f_{\pi}$ are the pion mass and decay constant, respectively. In particular, the axion mass (at a given $f_{a}$) and thus the critical density are suppressed by a factor of roughly $z^{\mathcal{N}}\simeq(1/2)^{\mathcal{N}}$ with respect to the standard QCD-axion. As shown in Balkin et al. (2024), values of $\mathcal{N}\geq 33$ lead to a new ground state of matter in white dwarfs and come with a large gap in the predicted radii of white dwarfs, which is incompatible with observations. For $\mathcal{N}=29$ and larger, the critical density is high enough for the metastable branch to explain all observed white dwarfs, and no exclusion can be placed. This leaves $\mathcal{N}=31$, which we focus on in this work.
@@
From the behaviour of the potential, we can now understand the behaviour of the field values and consequently the EOS. Starting from the unsourced case, the axion field sits at 0. Due to the potential barrier, the axion field remains in the metastable minimum at 0 until it fully vanishes at the critical density. This leads to a metastable branch of the EOS and is shown in red in Fig. 10. At densities above the critical density, the axion field can simply track the global minimum of the finite density effective potential at $a=\pi f_{a}$. If the axion has been sourced once within a star, it will continue to track the global minimum of the potential even at densities below $n_{c}$. For densities below the critical density, this global minimum at first stays at $\pi f_{a}$ and then continuously shifts towards the zero-density minimum closest to $\pi f_{a}$, at $\pi\left(1-\frac{1}{\mathcal{N}}\right)f_{a}$ as marked by the arrows in Fig. 10. Only at (near) zero-density, when the energy gain from the potential becomes lower than the gradient energy required for the transition to the true minimum, does the field return to the outside value of $a=0$.